Topic: Normal approximation to binomial distribution

Calculate the following probabilities using a normal approximation. P(9 ≤ X ≤ 12) where X ∼ B(21, 0.5)

Please show work as I will studying it step by step, thanks.
Answer #1

solution

given

X ∼ B(21, 0.5)

mean=n*p= 21*0.5=10.5

standard deviation = sqrt(n*p*(1-p))

=sqrt(21*0.5*(1-0.5)) =2.29128

SO P(9<=X<=12) = P((9-10.5)/2.29128 <(X-mean)/s <(12-.10.5)/2.29128)

=P(-0.6546 <Z< 0.6546)

=0.4844 (from standard normal table)